thie equation is identity, conditional,contradiction equation
4(x+7)=2(x+12)+2(x+1)

Answer by ewatrrr(24785)

4(x+7)=2(x+12)+2(x+1)
4x + 28 = 2x + 24 + 2x + 2
4x + 28 = 4x + 26
Contradiction